OpenStudy (anonymous):

please help help help with a couple.. medal to it A delivery truck is transporting boxes of two sizes: large and small. The combined weight of a large box and a small box is 70 pounds. The truck is transporting 70 large boxes and 65 small boxes. If the truck is carrying a total of 4775 pounds in boxes, how much does each type of box weigh?

OpenStudy (anonymous):

large box= small box=

OpenStudy (amoodarya):

large box= x small box=y x+y=70 70x+65y=4775 now you would be able to solve the system of equation ?

OpenStudy (anonymous):

A family has two cars. The first car has a fuel efficiency of 30 miles per gallon of gas and the second has a fuel efficiency of 20 miles per gallon of gas. During one particular week, the two cars went a combined total of 1600 miles, for a total gas consumption of 60 gallons. How many gallons were consumed by each of the two cars that week?

OpenStudy (anonymous):

1st car 2nd car

OpenStudy (anonymous):

@amoodarya

OpenStudy (anonymous):

howd you do it?

OpenStudy (amoodarya):

30x+20y=1600 x+y=60

OpenStudy (anonymous):

So, am I dividing? @amoodarya ?

OpenStudy (anonymous):

So, am I dividing? @amoodarya ?

OpenStudy (amoodarya):

dividing by ?

OpenStudy (amoodarya):

you need to solve the system by (any) method 30x+20y=1600 /10 --->3x+2y=160 x+y=60

OpenStudy (amoodarya):

for example x=60-y 3x+2y=160 3(60-y)+2y=160 180-3y+2y=160 y=20 ---> x=60-20=40